### v1

# Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Control Framework

Abstract
In the emerging hybrid traffic flow environment, which includes both human-driven vehicles (HDVs) and autonomous vehicles (AVs), ensuring safe and robust decision-making and control is crucial for the effective operation of autonomous vehicle platooning. Current systems for cooperative adaptive cruise control and lane changing are inadequate in responding to real-world emergency situations, limiting the potential of autonomous vehicle platooning technology. To address the aforementioned challenges, we propose a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven autonomous vehicle platooning Cooperative Control Framework. Within this framework, a deep reinforcement learning formation decision model integrating traffic priors is designed, and a twin-world deduction model based on safety priority judgment is proposed. Subsequently, an optimal control-based multi-scenario decision-control right adaptive switching mechanism is designed to achieve adaptive switching between data-driven and model-driven methods. Through simulation experiments and hardware-in-loop tests, our algorithm has demonstrated excellent performance in terms of safety, robustness, and flexibility. A detailed account of the validation results for the model can be found in Our Website.
Index Terms:
Autonomous Vehicle Platooning; Deep Reinforcement Learning; Mixed Traffic Flows; Self-Organised Cooperative Control.
I Introduction
The advent of autonomous driving and vehicle communication technologies has sparked significant interest in connected self-driving vehicles. Among these, autonomous vehicle platooning stands out for its potential to enhance traffic efficiency and safety [1, 2] in a relatively straightforward and cost-effective manner. By reducing aerodynamic drag, vehicle platooning can lower fuel consumption and emissions by at least 10%[3, 4],thereby alleviating environmental pollution[5, 6], while also decreasing road space usage and traffic congestion[7].
Most existing studies on autonomous vehicle platooning primarily focus on cooperative adaptive cruise control (CACC) and the cooperative lane change control (CLCC) to achieve a narrower inter-vehicle gap and a lower fuel consumption. Current autonomous driving platoons generally use LiDAR for virtual connectivity and communicate via wireless communication technologies[8]. Different communication protocols, including DSCR[9], GPSR[10], and GAHG[11], are utilized in various platooning algorithms. Different platoon structures have also been explored, which can be categorized into fully autonomous platoons and heterogeneous platoons with partially HDVs. The heterogeneous platoons can be further divided into those where the lead vehicle is human-driven and those where both the lead and tail vehicles are human-driven. Given the future development trends, this study primarily investigates control strategies for fully autonomous vehicle platoons.
CACC systems are integral for maintaining reduced inter-vehicle gaps and achieving lower fuel consumption, with a primary emphasis on ensuring the string stability[12, 13] of the platoon. These systems can be classified into model-based and rule-based methods. Model-based approaches, such as Model Predictive Control (MPC)[14, 15, 16, 17], leverage vehicle dynamics models and optimization techniques to achieve precise control of acceleration and deceleration[18]. However, they are often computationally demanding and may not adapt effectively to rapidly changing conditions. In contrast, rule-based methods rely on predefined rules and logic to dictate vehicle behavior[19, 20, 21], offering simplicity and faster computation at the cost of reduced flexibility in complex environments. In addition to these traditional methods, data-driven approaches, particularly those based on Reinforcement Learning (RL), have emerged as a promising alternative[13]. RL-based CACC systems can optimize control policies by learning from real-time data[22], effectively managing complex dynamics and improving response to changing traffic conditions. Despite their potential, these methods require substantial training data and pose challenges in ensuring safety during the learning phase.
CLCC enhances the maneuverability and efficiency of vehicle platoons by enabling coordinated lane changes. Advanced CLCC strategies, similar to CACC, include both model-based and rule-based approaches[23, 24, 25]. Model-based methods involve predictive control techniques to optimize lane change maneuvers based on anticipated future vehicle states[26, 27], while rule-based methods utilize a set of established rules to manage lane changes according to current traffic conditions[28, 29, 30]. Nevertheless, both approaches face limitations; model-based methods can be computationally intensive and sensitive to model inaccuracies, while rule-based methods may lack the adaptability required for managing complex traffic scenarios involving multiple interacting vehicles. RL-based approaches in CLCC similarly offer greater adaptability by learning from diverse traffic scenarios, allowing for optimized lane-changing strategies[31]. While these methods enhance flexibility, they require significant data and computational resources, and ensuring reliability in mixed traffic environments remains a challenge.
Current CACC and CLCC methods have notable shortcomings, particularly in handling scenarios with multiple types of strong interference. Existing strategies are typically designed to cope with one or two types of disturbances, such as sensor noise or communication delays, but struggle when faced with more complex interference patterns [32]. This limitation can lead to safety-critical situations, especially in mixed traffic environments where unpredictable HDVs are present. Consequently, the current state of autonomous vehicle platooning remains relatively conservative, often prioritizing safety over efficiency to mitigate risks.
To address these limitations, insights can be drawn from the robotics and Unmanned Aerial Vehicle (UAV) domain, where emergency obstacle avoidance and platoon reconfiguration strategies have been actively researched [33, 34, 35]. In robotics, these strategies often employ techniques such as real-time path planning[36, 37], dynamic reconfiguration algorithms, and machine learning-based decision-making to ensure robust performance in uncertain environments. These methods have demonstrated success in controlled and relatively closed environments, such as factory floors or isolated test tracks.
However, the application of these advanced strategies to autonomous vehicle platoons is limited by the open and highly variable nature of mixed traffic environments. Unlike controlled environments, mixed traffic flows are subject to unpredictable human behavior and diverse vehicle types, which impose significant challenges on the direct transfer of robotic strategies to autonomous driving. Despite these challenges, adapting and refining these approaches for use in vehicle platooning could potentially enhance the robustness and safety of autonomous driving systems.
In light of the aforementioned challenges, this study proposes a novel approach to autonomous vehicle platooning control. We propose a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework. This framework achieves adaptive decision-making for platooning in various complex scenarios by flexibly switching between deep reinforcement learning (DRL) integrated with traffic prior knowledge and model control methods. Firstly, the vehicle platooning problem is modeled as a multi-agent cooperative decision-making problem, and a RL decision-making framework based on Proximal Policy Optimization (PPO)[36] is proposed, with a decision network structure and reward function designed based on traffic prior knowledge. A twin-world model based on safety priority judgment is introduced and coupled with the DRL algorithm to enhance algorithm safety. Subsequently, combining optimal control methods, a multi-scenario decision-control right adaptive switching mechanism is designed. To evaluate our method, we have devised intricate and high-risk mixed-traffic scenarios, as depicted in Figure 1. These scenarios encompass various elements such as human interference, traffic accidents, and fluctuations in traffic flow. Through simulations and hardware-in-the-loop experiments, our algorithm has been validated in terms of safety, robustness, and flexibility.
Figure 1: Illustration of the considered traffic scenario and the solution. CAVs (Red) and HDVs (Green) are both considered
Our contributions are summarized as follows:
  • 1)
    A Data-Model-Knowledeg Hybrid-Driven Framework for autonomous vehicle platooning is constructed, which integrates data-driven methods, model-based control, and domain-specific knowledge, enabling autonomous vehicle platoons to achieve adaptive decision-making and robust control in complex and dynamic traffic environments.
  • 2)
    A twin-world safety enhancement model is coupled with RL and implement an adaptive switching strategy. The mechanism enables the platoon to seamlessly switch between routine and safety-critical operations.
  • 3)
    A Cyber-Physical Experimentation Platform is established, incorporating simulation and real-world controllers, to validate the framework’s robustness and feasibility under diverse environmental conditions.
To better present our work, the rest of this paper is organized as follows. Section II presents the hybrid-driven self-organizing control framework. Section III introduces in detail the proposed Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework Section IV presents the performance evaluation. Finally, we conclude the paper in Section V.
II Problem statement and framework review
II-A Scenario Description
As the most common application scenario for autonomous vehicle platooning, we have conducted an in-depth study on the self-organizing decision-making problem of multi-vehicle platooning on a three-lane highway with strong disturbances. Based on real-world issues encountered in various platooning tests, we identified and summarized several major types of strong disturbances, including human interference, traffic accidents, and traffic flow oscillations. In order to address these disturbances, a series of representative scenarios were devised, including emergency braking of the leading vehicle, sudden traffic accidents, and medium to low-speed traffic flow scenarios. These are illustrated in Figure 1.
Conventional obstacle avoidance strategies for autonomous vehicle platooning often struggle to handle these strong disturbance factors. For the practical implementation of autonomous vehicle platooning, it is essential to safely navigate through regions with strong disturbances and ultimately maintain stable operation at high speeds to reach the destination.
In this study, the high-level action decisions for the autonomous vehicle platoon are determined by our proposed Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven framework. This algorithm then utilizes a closed-loop Proportional-Integral-Derivative (PID) controller to translate these decisions into low-level steering and acceleration/deceleration controls.
II-B Framework Overview
Figure 2: The overview of the proposed data-model-knowledge hybrid-driven framework for self-organizing autonomous vehicle platooning
To address emergency risk scenarios that traditional vehicle platooning technologies cannot handle, this study proposes a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework. As illustrated in Figure 2, the framework is structured into two primary layers:
  • Upper Layer: This layer utilizes a Proximal Policy Optimization DRL decision model following the Centralized Training and Decentralized Execution (CTDE) paradigm. The DRL model is coupled with a twin-world model to perform safety verifications and ensure the generation of safe decisions.
  • Lower Layer:This layer implements a Linear-Quadratic Regulator (LQR) for the longitudinal control of platoon vehicles, aimed at maintaining smaller inter-vehicle headway distances.
These decision-making strategies are adaptively and flexibly switched using a knowledge-based finite state machine. The LQR controller is employed to maintain smaller headway distances in safe driving environments, thereby enhancing the platoon’s efficiency. In complex and challenging scenarios, the reinforcement learning model assumes control, managing the splitting and reformation of the platoon to ensure more flexible and safe driving.
III Methodology
This section details the algorithm design for autonomous vehicle platoon reconfiguration and maintenance. First, the problem is modeled as a Partially Observable Markov Decision Process (POMDP)[38], with a clear explanation of the observation space, action space, and reward function, integrating prior traffic knowledge. Then, the safety-priority-based twin-world model is introduced. Finally, this study combines optimal control methods to design an adaptive switching mechanism suited for multi-scenario applications, referred to as the platoon maintenance model.
III-A Problem Formulation and Reinforcement Learning Setup
Driving platoon vehicles on highways is a complex and challenging task, especially in mixed traffic flows. Treating the platoon as a single entity can lead to safety risks under strong disturbances, requiring reconfiguration. This process involves multi-agent cooperative decision-making[39], as each vehicle’s state is influenced by the actions of other platoon members and surrounding traffic.
Traditional decision algorithms are inadequate for this task; thus, the problem should be modeled as a POMDP to employ deep reinforcement learning. The challenge lies in balancing individual and collective goals while navigating a dynamic environment where actions influence both nearby and distant vehicles. The system’s schematics are shown in Figure 3.
In this study, we model the problem of multi-vehicle platooning on highways as an agent that accesses an observation at each time $t$. Given the sensor capabilities of platoon vehicles and the complexity of real traffic environments, the observation only provides partial information about the system state $s_{t}$. The platoon cannot accurately perceive the actions and intentions of all other vehicles, making the observation space $o_{t}$ a proper subset of the system state $s_{t}$.
Using $o_{t}$, the platoon can sample decision actions $a_{t}$ from the action space according to the shared decision policy $\pi_{\theta}$, as follows:
$a_{t}\sim\pi_{\theta}\left(a_{t}\mid o_{t}\right)$ (1)
where the variable $\theta$ represents the trained policy parameters. The agents within the platoon execute action $a_{t}$, interact with the environment, and transition the system to state $s_{t+1}$, acquiring a new observation $o_{t+1}$.
The vehicle platoon, while making decisions as a whole, employs a decentralized execution strategy. Therefore, we can formulate this problem as a POMDP, which can be described by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{O})$, where $\mathcal{S}$ is the set of system states, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the state transition model, $\mathcal{R}$ is the reward function, and $\mathcal{O}$ is the observation space. The observation space $\mathcal{O}$ is determined by the system state and the position of the platoon $({o_{t}}\in{s_{t}})$.
In the POMDP, the agent makes decisions based on the policy $\pi$ and the current observation space. The objective of the agent is to find the optimal policy $\pi^{*}$ that maximizes the cumulative future rewards[40]. The optimal action-value function $Q^{\pi}(s,a)$, which can be defined as follows:
$Q^{\pi}(s,a)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$ (2)
The subsequent section will present the design of the observation space $\mathcal{O}$, action space $\mathcal{A}$, and reward function $\mathcal{R}$.
  • Action space: In this study, each agent $\text{agent}_{i}$ has an action space that includes a set of high-level control decisions: left turn, right turn, accelerate, decelerate, and cruise[41, 42]. For vehicle platooning, we integrate the action spaces of multiple agents using a quinary encoding scheme. This approach encodes the combined action space using a shared encoder, forming a 125-dimensional discrete action space through the Cartesian product of the individual actions. The overall action space for the platoon is thus the joint action space of all autonomous vehicles, denoted as $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$.
    Figure 3: The schematics of the system and the simulation setup
    As illustrated in Figure 3, after making high-level decisions, the platoon’s bottom-level controllers manage each vehicle’s steering and throttle, generating specific control signals to execute vehicle movements in the simulation environment.
  • Observation Space: The observation space can be represented as a matrix with dimensions $|N_{i}|\times|F|$, where $|N_{i}|$ is the number of vehicles observed by the agent, and $|F|$ represents the feature vector of other vehicles’ positions and actions. The feature vector can be defined as:
    $F_{i}=[x_{i},y_{i},v_{i}^{x},v_{i}^{y},is\_in\_platoon]$ (3)
    To better represent the interactions between the platoon and surrounding vehicles, $x_{i},y_{i},v_{i}^{x},v_{i}^{y}$ are the relative positions and velocities of other vehicles with respect to the observing vehicle, indicating longitudinal position, lateral position, longitudinal velocity, and lateral velocity, respectively. The variable ${is\_in\_platoon}$ is a binary indicator (0 or 1) that denotes whether the observed vehicle is part of the platoon.
    In this study, we consider the sensor detection range limitations and assume that only nearby vehicles can be observed. Specifically, only vehicles within $D_{vision}$ ahead or behind the platoon can be detected[43]. The observation space for the entire platoon system is the Cartesian product of the individual vehicles’ observation spaces, denoted as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\times\mathcal{S}_{N}$.
  • Reward Function: The reward function is a crucial component in training the DRL model, as it determines the model’s overall performance. This study designs both single-vehicle rewards and multi-vehicle collaborative rewards to ensure the robustness and safety of the platoon’s decisions. The overall reward for the vehicle platoon is the sum of the rewards obtained by each vehicle and the collective rewards:
    $R_{global}(s,a)=\frac{r_{L}}{n}\sum_{i=1}^{n}R_{ind}^{i}+R_{sys}$ (4)
    $R_{ind}(s,a)=\sum_{t\in\{C,F,A\}}w_{t}\times r_{t}$ (5)
    $R_{sys}(s,a)=\sum_{k\in\{M,D,H,S\}}w_{k}\times r_{k}$ (6)
    where $w_{C}$, $w_{L}$, $w_{F}$, and $w_{A}$ are the weights for the single-vehicle rewards: collision reward $r_{C}$, lane reward $r_{L}$, high-speed reward $r_{F}$, and acceleration reward $r_{A}$, respectively. Since collisions are unacceptable, the absolute value of $w_{C}$ will be significantly greater than the other coefficients. $w_{M}$, $w_{D}$, $w_{H}$, and $w_{S}$ correspond to the multi-vehicle collaborative rewards: same-lane reward $r_{M}$, distance traveled reward $r_{D}$, headway reward $r_{H}$, and speed coordination reward $r_{S}$. Single-vehicle rewards primarily ensure the safety and efficiency of individual vehicles, while multi-vehicle rewards maintain the overall structure of the platoon, preventing disbandment issues. We considere both individual vehicle rewards and multi-vehicle collaborative rewards. For individual vehicle rewards, the focus is on safety and efficiency. In contrast, multi-vehicle collaborative rewards emphasize the cooperation and coordination among vehicles. The schematic diagram of the reward function is shown in Figure 8.
    Figure 4: Illustration of the designed headway maintaining reward, which considers both the single-vehicle rewards and the multi-vehicle rewards
III-B Deep Reinforcement Learning Algorithm
Traditional DRL algorithms such as Deep Q-Networks (DQN) have been applied to autonomous driving decision-making[44]. However, these algorithms primarily focus on single-agent learning and perform inadequately in scenarios requiring complex multi-vehicle cooperative decision-making. This study addresses the challenge of vehicle platooning reconstruction in mixed traffic flow, aiming to develop a robust self-organizing decision model for vehicle platooning under highly disruptive conditions through agent collaboration.
To this end, we have adapted the PPO algorithm[45] to enhance its application to the safe and robust multi-vehicle cooperative decision-making problem. PPO is distinguished by its balance of simplicity, sample efficiency, and robustness, and has shown exceptional performance in various tasks.
In our approach, we utilize a CTDE strategy. As illustrated in Figure 5, the sensor data from all platooning vehicles are aggregated into a unified observation space. A shared policy $\pi_{\theta}$ then outputs the actions for all vehicles within the platoon. Following the execution of these actions by the low-level controllers, the agents update the policy $\pi_{\theta}$ based on rewards obtained from environmental feedback, employing experience replay and gradient-based optimization.
Figure 5: The illustration of the PPO-based formation configuration algorithm schematic model
In this self-organizing autonomous vehicle platooning system, all agents are treated equally and receive the same global reward. The objective of the DRL optimization is to maximize the global reward, defined as:
$\max{R_{g,t}=\sum_{i=1}^{N}r_{g,t}}$ (7)
In deep reinforcement learning, deep neural networks (DNNs) are used to approximate policies, enabling agents to learn how to make decisions within our environment to maximize cumulative rewards. DNNs play a crucial role in decision-making within deep reinforcement learning. This study employs the Actor-Critic architecture-based DRL neural network structure depicted in Figure 6.
Figure 6: Schematic diagram of the network structure based on the Actor-Critic architecture
In this architecture, the policy loss and value function error loss are combined into a single loss function. Under the condition of shared parameters among agents, the total loss function can be expressed as a linear combination of the policy loss, state update loss, and entropy regularization term[45]:
$J\left(\theta_{i}\right)=J^{\pi_{\theta_{i}}}-\beta_{1}J^{V_{\phi_{i}}}+\beta_{2}H\left(\pi_{\theta_{i}}\left(s_{t}\right)\right)$ (8)
where $\beta_{1}$ and $\beta_{2}$ are the weight values for the value loss function and the entropy regularization term, respectively. The policy loss $J^{\pi_{\theta_{i}}}$ and the state update loss function $J^{V_{\phi_{i}}}$ are defined as:
$J^{\pi_{\theta_{i}}}=E_{\pi_{\theta_{i}}}\left[\log\pi_{\theta_{i}}\left(a_{i,t}|s_{i,t}\right)A_{i,t}^{\pi_{\theta_{i}}}\right]$ (9)
$J^{V_{\phi_{i}}}=\min_{\phi_{i}}E_{\mathcal{D}_{i}}\left[\left(r_{i,t}+\gamma V_{\phi_{i}}\left(s_{i,t+1}\right)-V_{\phi_{i}}\left(s_{i,t}\right)\right)\right]^{2}$ (10)
where $A_{i,t}^{\pi_{\theta_{i}}}$ is the advantage function, defined as:
$A_{i,t}^{\pi_{\theta_{i}}}=r_{i,t}+\gamma V^{\pi_{\phi_{i}}}(s_{i,t+1})-V^{\pi_{\phi_{i}}}(s_{i,t})$ (11)
where $V^{\pi_{\phi_{i}}}(s_{i,t})$ is the state value function. Additionally, the entropy regularization term encourages the agent to explore new directions and avoid potential local optima[46]. It is expressed as:
$H\left(\pi_{\theta_{i}}\left(s_{t}\right)\right)=E_{\pi_{\theta_{i}}}\left[-\log\left(\pi_{\theta_{i}}\left(s_{t}\right)\right)\right]$ (12)
This study updates the network parameters using mini-batch optimization sampling. The optimization objective is to minimize the variance of Equation 11.
III-C Twin World Security Projection
While DRL enables robust decision-making in mixed traffic with strong interferences, its ”black-box” nature can lead to unexpected outcomes in standard scenarios and may not fully comply with traffic regulations. Although more reward functions can be used to impose these constraints, this approach can result in significant training costs and potential model divergence. This study addresses these issues by implementing a comprehensive and extensible safety mask.
To ensure decision safety, we propose a safety-priority-based twin-world model. This model evaluates the safety of DRL decisions by first detecting potential collisions based on safety priorities, thereby reducing computational complexity. It then recursively computes $T_{n}$ steps in a twin world constructed from the current scenario, optimizing actions based on the platoon’s safety margin.
Given the focus of this study on decision-making in mixed traffic, it is essential to predict the driving behavior of HDVs. To this end, we utilize the prevalent Intelligent Driver Model (IDM) [47] for simulating car-following behaviors. For lane-changing decisions, we adopt the MOBIL model [23]. These two models, when combined, ensure a minimal risk of collisions among HDVs while closely mirroring human driving patterns.
The safety mask models each vehicle as a circle with the vehicle’s center as the center and half the vehicle’s length as the radius. We simulate future $T_{n}$ steps using motion models of HDVs and CAVs to identify potential collision risks. While straightforward for single vehicles, this simulation becomes computationally challenging for vehicle platoons due to the exponentially increasing action space. To mitigate this, we propose a safety-priority-based simulation model. This model prioritizes vehicles with the smallest safety margins and conducts safety simulations in the twin world. The entire platoon tends to maintain cohesive formation, allowing for high computational efficiency suitable for real-time applications.
Safety priority is determined based on headway time, with higher priority assigned to vehicles with smaller headway times relative to HDVs. The safety priority $p_{i}$ is defined as:
$p_{i}=-\log\frac{d_{headway}}{v_{t}}+\sigma_{i}$ (13)
where $d_{headway}$ is the headway distance, $v_{t}$ is the current vehicle speed, and $\sigma_{i}$ is a small random term to avoid identical safety priorities.
During the safety simulation, the vehicle with the highest safety priority undergoes assessment at each time step $t$. The twin world checks whether the vehicle’s decisions over the next $T_{n}$ time steps conflict with nearby vehicles. If no collision is detected (i.e., the minimum distance between vehicles exceeds the threshold), the DRL decision is deemed safe. Otherwise, the twin world traverses other possible decisions, selecting the action that maximizes safety margins:
$a_{t}^{\prime}=\arg\max_{a_{t}\in\mathcal{A}_{\mathrm{pro}}}\left(\min_{k\in T_{n}}d_{\mathrm{sm},k}\right)$ (14)
where $\mathcal{A}_{\mathrm{pro}}$ represents alternative actions output by the DRL, and $d_{\mathrm{sm}}$ refers to the safety margin, defined as the minimum headway distance in both current and target lanes during a left or right turn. The detailed procedure is illustrated in Algorithm 1.
Input: $vehicles$,$x_{i}$,$y_{i}$,$v_{x}^{i}$,$v_{y}^{i}$,$headway_{i}$
Output: $a_{i,i\in formation}$
1
2for $i=0$ to $N$ do
3       Compute safety priority: $priority_{i}=-\ln\left({headway_{i}}/{v_{i}}\right)$;
4       Sort vehicles by safety priority and store in list $\phi$;
5
6 end for
7for $j=0$ to $|\phi|$ do
8       Retrieve the vehicle with the highest safety priority: $\phi[0]$;
9       Identify neighboring vehicles and store them in list $N_{\phi[0]}$;
10       Predict the trajectory for the next $T_{n}$ simulation steps: $\xi_{v},v\in\phi[0]\cup N_{\phi[0]}$;
11       if trajectories overlap then
12             Calculate safety margin $d_{\text{sm},i}$;
13             if lane changed then
14                   $d_{\text{sm},i}=\min_{j}\left(p_{i}-p_{j},j\in\text{objective}\,v_{i}\right)$
15             end if
16            else
17                   $d_{\text{sm},i}=\text{headway}\,v_{i}$
18             end if
19            Replace action $a_{t}\leftarrow a_{t}^{\prime}$;
20             Update trajectory list $\xi_{\phi[0]}\leftarrow\xi_{\phi[0]}^{\prime}$;
21
22       end if
23      Replace action $a_{t}\leftarrow a_{t}^{\prime}$ Update $\phi[i]\leftarrow\phi[i+1],i=1,2,\cdots$
24 end for
return $a_{i,i\in formation}$
Algorithm 1 Twin World Security Projection
III-D Formation maintenance
The platoon reconfiguration model ensures safe and robust self-organizing decision-making and control in complex scenarios with strong interference. However, it struggles to maintain small headway distances in routine scenarios due to overly conservative strategies aimed at avoiding collision risks. To address this issue, we propose an adaptive switching mechanism, the platoon maintenance model, which integrates optimal control methods.
This model maintains stable, close headway distances in routine scenarios while dynamically adjusting driving strategies based on traffic conditions and interference levels. By monitoring the environment and vehicle states in real-time, the model adapts its strategies to ensure safety in complex scenarios and efficiency in routine conditions. This approach enhances overall coordination and driving efficiency, providing a flexible response to varying conditions.
To address the issue of decision control allocation in vehicle platooning across different scenarios, this study employs a Finite State Machine (FSM) for decision-making. At any given time, the state automaton can be in only one state. The FSM is defined by the five-tuple $M(Q,D,\delta,q_{0},F)$, where:
  1. 1.
    $Q=\{S_{1},S_{2}\}$ is the set of states;
  2. 2.
    $q\in Q$ represents the current state;
  3. 3.
    $\delta$ is the state transition function;
  4. 4.
    $F$ is the set of output states
  5. 5.
    $D$ denotes the risk level of the traffic scenario
This study considers a scenario as a routine safe scenario when the platoon is on the same lane and there are no other traffic participants within $L_{safe}$ ahead or behind the platoon. In this case, the FSM switches to state $S_{1}$, adopting the LQR control model[48] to ensure smaller headway distances. In this study, string stability within the platoon[12] is not emphasized, as it has been thoroughly addressed in previous research. In all other scenarios, the FSM switches to state $S_{2}$, which employs a safety-enhanced data-driven strategy to balance the safety and efficiency of the platoon.
By defining these states and transitions, the FSM provides a structured approach to adaptively manage the control strategies of the platoon, ensuring optimal performance across varying traffic conditions. The detailed procedure is illustrated in Algorithm 2.
Input: $vehicles$, $riskLevel$, $Participants$, $L_{safe}$
Output: $controlStrategy$
1
2for $i=0$ to $N$ do
3       Assess risk level and number of traffic participants within $L_{safe}$;
4       if ($riskLevel>=RoutineSafe$) and (Participants within $L_{safe}$) then
5             Switch to State $S_{1}$: LQR Control Model;
6             Activate LQR Control;
7             Set $controlStrategy=LQR$;
8
9      else
10             Switch to State $S_{2}$: Safety-Enhanced Data-Driven Strategy;
11             Set $controlStrategy=DataDriven$;
12
13       end if
14      Execute $controlStrategy$ based on the current state;
15       if $controlStrategy==LQR$ then
16             ControlWithLQR();
17
18       end if
19      else if $controlStrategy==DataDriven$ then
20             ControlWithDataDrivenStrategy();
21
22       end if
23
24 end for
return $controlStrategy$
Algorithm 2 Adaptive Switching Mechanism
IV Simulation and Performance Evaluation
In this section, the software-in-the-loop (SIL) experiments are conducted to validate the effectiveness of the proposed method. Additionally, a hardware-in-the-loop (HIL) experimental platform is built to further verify the model’s safety and robustness and to assess its feasibility for direct deployment in real vehicles.
IV-A Software-in-the-loop experiment
IV-A1 Environment and simulation setting
The simulation environment is constructed upon the Highway-Env simulator[49], a reinforcement learning experimental environment library based on an OpenAI Gym environment[50]. As illustrated in Figure 3, all vehicles’ high-level policies are transformed into throttle and steering information through a low-level PID controller, thereby enabling vehicle movement within the simulation environment. Furthermore, our HDVs employ the IDM[47] car-following model and the MOBIL[23] lane-changing model provided by Highway-Env to achieve effective control in both longitudinal and lateral directions.
During the training process, a varying number of HDVs will randomly appear at randomly generated refresh points. The control sampling frequency is set to $15$ Hz, meaning decisions and controls are made 15 times per second.
We use three different random seeds and trained the proposed model in various complex scenarios, with each training session running for over 100,000 steps until the model converged. We selecte the Deep Q-Learning (DQN)[51] and Advantage Actor-Critic (A2C)[52] algorithms as experimental baselines. The training parameters for the DRL model are detailed in Table I.
TABLE I: Hyperparameter settings for reinforcement learning
Param Definition Value
$\eta$ Learning Rate $5\times{10}^{-4}$
$N_{steps}$ steps for each environment per update $256$
$N_{t}$ Total Training Steps $10^{5}$
$\beta_{1}$ Loss function weights 1
$\beta_{2}$ Loss function weights 0.01
$\lambda$ Minibatch size 128
$\gamma$ Discount factor 0.8
$w_{C}$ Weight for $r_{C}$200
$w_{F}$ Weight for $r_{F}$ 5
$w_{A}$ Weight for $r_{A}$ 1
$w_{M}$ Weight for $r_{M}$ 1.2
$w_{D}$ Weight for $r_{D}$ 1.2
$w_{H}$ Weight for $r_{H}$ 1
$w_{S}$ Weight for $r_{S}$ 2
IV-A2 Performance evaluation
In this subsection, we evaluate the performance of the proposed model within a simulated training environment, comparing it with advanced RL algorithms (e.g., DQN, A2C) and traditional CACC and CLCC platooning strategies. Our results indicate that the proposed vehicle platooning model significantly outperforms existing models across various scenarios.
We designed the scenario depicted in Figure 7 not merely as a training environment but as an edge case that stresses the safety and robustness of the model under extreme conditions. The platoon vehicles travel at a speed of 28 m/s within a mixed traffic flow, where surrounding vehicles’ speed distribution follows a uniform range of 25 to 28 m/s. The traffic flow is characterized by oscillations and diverse driving styles, including aggressive, conservative, and neutral behaviors, creating a volatile environment prone to accidents. Furthermore, we introduced a few malfunctioning vehicles that may suddenly stall or brake unexpectedly, further heightening the risk. The average inter-vehicle distance for HDVs is approximately 20 meters, providing insufficient space for the platoon to execute coordinated lane changes. Our study trained the model in this complex and perilous scenario, incorporating randomly generated traffic flows to simulate real-world conditions.
Figure 7: Schematic representation of the model training scenario wherein a multitude of interferences are present
  1. 1.
    Sample Efficiency: As shown in Figure 8, our safety-enhanced RL algorithm exhibits superior sample efficiency compared to baseline algorithms. The model achieves higher rewards per time step, rapidly increasing reward values within approximately 20,000 time steps and reaching initial convergence around 40,000 time steps. Despite typical training fluctuations, the model maintains an upward trend and stabilizes, indicating successful convergence. The consistently higher rewards across different scenarios are due to the model’s early prevention of illogical and illegal actions, reducing exploration time and enhancing operational efficiency.
    TABLE II: Testing performance comparison of the collision rate, average speed and average vehicle spacing
    of proposed model and a conventional vehicle formation model.
    Baseline DRL Model DRL + Formation Maintenance Proposed Model
    Collision rate(%) 61.25 5.03 5.40 3.00
    Average Speed(km/h) 59.09 73.87 74.08 74.98
    Average spacing(m) 5.92 5.17 3.10 3.01
    Figure 8: Schematic representation of the model training reward function in multiple scenarios.
  2. 2.
    Scenario Evaluation: We focused on traffic flow oscillations, a common and representative scenario, to assess the performance of the traditional CACC and CLCC, a reinforcement learning-based model, and our proposed model. We analyzed the time-varying behavior of headway, acceleration, speed, and lateral position under these conditions, as illustrated in Figure 9.
    Figure 9: Schematic illustration of headway, acceleration, velocity and lateral position of vehicle formations in three typical risk cases.
    Safety: The traditional CACC and CLCC, influenced by the IDM-lead vehicle, displayed significant sensitivity to the motion of preceding vehicles, leading to frequent adjustments in acceleration and speed. This resulted in unstable behavior, particularly during traffic flow oscillations. In contrast, both the reinforcement learning-based model and our proposed model executed adaptive lane changes by the 7th time step, effectively mitigating the impact of oscillations and enhancing overall safety.
    Efficiency: The traditional CACC and CLCC, lacking platoon reconfiguration capabilities, achieved efficiency gains only in the presence of large gaps. Concurrently, the intricate traffic patterns have precluded CLCC from identifying viable locations for collaborative lane changes, thereby impeding the potential for enhanced operational efficiency. Despite maintaining smaller headways, its lower speed reduced overall efficiency. The reinforcement learning-based model improved platoon speed by reconfiguring to more optimal lane positions. Our proposed model further enhanced efficiency by maintaining the platoon post-reconfiguration, sustaining high speeds and smaller headways, leading to more stable and efficient operation.
    To generalize our findings, we tested 30,000 randomly generated high-interference scenarios to evaluate three models: the DRL model, the DRL model with formation-keeping, and our complete proposed model, including the twin-world model. The baseline platoon consisted of an IDM lead vehicle with LQR-following vehicles maintaining a 0.5-second headway.
    In contrast, traditional CACC and CLCC models exhibit high collision rates in high-interference scenarios. However, our proposed method significantly reduces these rates while also enhancing efficiency. It is important to note that due to the randomness and complexity of generated scenarios, some unavoidable situations may still lead to collisions across all algorithms. As shown in Table II, our proposed model outperforms traditional strategies across multiple metrics. The DRL model effectively handles emergency scenarios, reducing the collision rate to below 5%. Incorporating formation-keeping maintains a comparable collision rate of 5.40%. However, the twin-world safety model further reduces the collision rate to 3%, underscoring the substantial impact of safety simulations.
    Moreover, the proposed model achieves an average platoon passing speed exceeding 70 km/h, a significant improvement over the baseline’s 59.09 km/h, indicating enhanced safety and optimized traffic flow. Additionally, the model supports smaller headways, achieving an average spacing of 3 meters compared to 5.92 meters with traditional methods, reflecting improved vehicle control and coordination.
    Overall, the proposed model outperforms traditional platooning strategies, demonstrating substantial improvements in safety, robustness, and efficiency.
IV-B Hardware-in-loop experiment
(a) HIL experimental platform framework
(b) HIL experiment platform physical diagram
Figure 10: Overall Architecture of the HIL Experimental Platform
Figure 11: Schematic representation of case 1 results in HIL experiments
Figure 12: Schematic representation of case 2 results in HIL experiments
IV-B1 HIL experimental platform construction
In the Hardware-in-the-Loop (HIL) experiments, we constructe an advanced simulation platform integrating a domain controller with Carla[53], a high-fidelity driving simulator, to validate our proposed model in a controlled yet realistic environment. The experimental highway scenario is created using Roadrunner, with varied background traffic flows generated using Highway-env.
TABLE III: Specifications of the domain controller
Parameter Specification
CPU Cortex-A78AE – Arm®
CPU Frequency 2.2 GHz
Memory 64 GB 256-bit 204.8 GB/s
GPU Performance 275 TOPS
Storage 2 TB SSD
The platform consists of the domain controller, Carla simulation environment, Roadrunner, Highway-env, and PID controllers. The domain controller, whose parameters are shown in Table III, processes real-time traffic and road information from Carla. Carla provides detailed road conditions and vehicle dynamics, which are translated into vehicle movements via PID controllers based on high-level action commands. Roadrunner generates complex highway scenarios, while Highway-env produces traffic with varied driving styles to test the algorithm’s robustness.
During experiments, the domain controller gathers real-time data from Carla, including vehicle positions, velocities, and traffic conditions.Based on this data, the proposed DRL algorithm makes high-level decisions, such as speed adjustments and lane changes. These decisions are sent back to Carla, where PID controllers convert them into specific throttle and steering commands. Various scenarios, including high-risk and normal conditions, are simulated to evaluate the algorithm’s performance based on metrics like safety, robustness, and efficiency. Figure 10 illustrates the platform’s overall architecture, highlighting its key components.
IV-B2 Case analysis
This section examines the performance of our model on the HIL simulation platform in real-time autonomous platooning scenarios. The results indicate that our model can effectively perform perception, decision-making, and control, ensuring safe and robust autonomous platooning. We analyze two typical scenarios involving significant static and dynamic interferences to illustrate the model’s strategies under these conditions. Demonstration videos are available on our website  [^1] Hardware-in-the-Loop Experimental Validation Video Weblink.
In the first scenario, the platoon encounters a severe static obstruction: a multi-vehicle collision blocking multiple lanes. Traditional CACC strategies cannot manage coordinated deceleration due to inadequate braking distances, and CLCC strategies are hindered by vehicles behind the accident on the right. Our model resolves this complex situation through adaptive decision-making, avoiding secondary collisions by allowing the first vehicle to initiate a lane change and the other vehicles to follow once adequate space is available.
As shown in Figure 11, the accident is detected at step 30. The first vehicle moves to the rightmost lane, completing the maneuver by step 105. By step 160, the remaining vehicles have sufficient space to change lanes, avoiding the collision zone, and by step 220, all vehicles have successfully moved. The LQR controller then stabilizes the platoon with reduced headway distances, demonstrating the model’s flexibility and efficiency in navigating static obstructions.
In the second scenario, the platoon faces a dynamic obstruction: the leading vehicle decelerates suddenly at 5 m/s² due to a breakdown. Traditional strategies would struggle to avoid collisions due to short following distances and the presence of background traffic. Our model, however, initiates a staggered lane change, as illustrated in Figure 12. The first vehicle moves left at step 20, allowing the second to move right. The third vehicle, facing traffic in the right lane, moves left. By step 70, all vehicles have repositioned and resumed normal speeds. The rightmost lane, offering higher efficiency, sees gradual reformation of the platoon, with all vehicles settled by step 455.
These scenarios demonstrate the model’s capability to handle sudden obstacles by dynamically adjusting the platoon’s formation and leveraging available lane space to maintain safety and efficiency. The adaptive lane-changing strategy ensures minimal disruption and optimal traffic flow, showcasing the model’s robustness and flexibility in complex driving conditions.
V Conclusion
In the complex environment of mixed human and autonomous driving, achieving robust and safe self-organizing decision-making and control for autonomous vehicle platooning remains a significant challenge for commercial deployment. This paper presents a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework for autonomous vehicle platoons. This framework enables adaptive decision-making in various complex scenarios by flexibly switching between reinforcement learning and model control methods, integrating traffic priors and optimal control techniques. Our model ensures safe and stable operation even under severe dynamic and static disturbances. Extensive simulations and HIL tests have demonstrated the effectiveness of the proposed model. The results show that our model significantly outperforms traditional platooning strategies such as CACC and CLCC in terms of safety, efficiency and robustness, demonstrating its potential for real-world deployment.
Overall, this research contributes to the advancement of autonomous vehicle platooning by providing a model that balances safety, efficiency, and robustness. Future work will focus on optimizing the model’s performance, exploring its scalability to larger platoons, and conducting real-world testing to confirm its practical benefits and limitations.

### v2

@@
-Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Control Framework
+Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Decision-making Framework
 Abstract
-In the emerging hybrid traffic flow environment, which includes both human-driven vehicles (HDVs) and autonomous vehicles (AVs), ensuring safe and robust decision-making and control is crucial for the effective operation of autonomous vehicle platooning.
-Current systems for cooperative adaptive cruise control and lane changing are inadequate in responding to real-world emergency situations, limiting the potential of autonomous vehicle platooning technology.
-To address the aforementioned challenges, we propose a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven autonomous vehicle platooning Cooperative Control Framework.
-Within this framework, a deep reinforcement learning formation decision model integrating traffic priors is designed, and a twin-world deduction model based on safety priority judgment is proposed.
-Subsequently, an optimal control-based multi-scenario decision-control right adaptive switching mechanism is designed to achieve adaptive switching between data-driven and model-driven methods.
-Through simulation experiments and hardware-in-loop tests, our algorithm has demonstrated excellent performance in terms of safety, robustness, and flexibility.
+In hybrid traffic environments where human-driven vehicles (HDVs) and autonomous vehicles (AVs) coexist, achieving safe and robust decision-making for AV platooning remains a complex challenge.
+Existing platooning systems often struggle with dynamic formation management and adaptability, especially in unpredictable, mixed-traffic conditions.
+To enhance autonomous vehicle platooning within these hybrid environments, this paper presents TriCoD, a twin-world safety-enhanced Data-Model-Knowledge Triple-Driven Cooperative Decision-making Framework.
+This framework integrates deep reinforcement learning (DRL) with model-driven approaches, enabling dynamic formation dissolution and reconfiguration through a safety-prioritized twin-world deduction mechanism.
+The DRL component augments traditional model-driven methods, enhancing both safety and operational efficiency, especially under emergency conditions.
+Additionally, an adaptive switching mechanism allows the system to seamlessly shift between data-driven and model-driven strategies based on real-time traffic demands, thereby optimizing decision-making ability and adaptability.
+Simulation experiments and hardware-in-the-loop tests demonstrate that the proposed framework significantly improves safety, robustness, and flexibility.
 A detailed account of the validation results for the model can be found in Our Website.
 Index Terms:
-Autonomous Vehicle Platooning; Deep Reinforcement Learning; Mixed Traffic Flows; Self-Organised Cooperative Control.
+Autonomous Vehicle Platooning; Deep Reinforcement Learning; Mixed Traffic Flows; Self-Organised Cooperative Decision-making.
 I Introduction
-The advent of autonomous driving and vehicle communication technologies has sparked significant interest in connected self-driving vehicles.
-Among these, autonomous vehicle platooning stands out for its potential to enhance traffic efficiency and safety [1, 2] in a relatively straightforward and cost-effective manner.
-By reducing aerodynamic drag, vehicle platooning can lower fuel consumption and emissions by at least 10%[3, 4],thereby alleviating environmental pollution[5, 6], while also decreasing road space usage and traffic congestion[7].
-Most existing studies on autonomous vehicle platooning primarily focus on cooperative adaptive cruise control (CACC) and the cooperative lane change control (CLCC) to achieve a narrower inter-vehicle gap and a lower fuel consumption.
-Current autonomous driving platoons generally use LiDAR for virtual connectivity and communicate via wireless communication technologies[8].
-Different communication protocols, including DSCR[9], GPSR[10], and GAHG[11], are utilized in various platooning algorithms.
-Different platoon structures have also been explored, which can be categorized into fully autonomous platoons and heterogeneous platoons with partially HDVs.
-The heterogeneous platoons can be further divided into those where the lead vehicle is human-driven and those where both the lead and tail vehicles are human-driven.
-Given the future development trends, this study primarily investigates control strategies for fully autonomous vehicle platoons.
-CACC systems are integral for maintaining reduced inter-vehicle gaps and achieving lower fuel consumption, with a primary emphasis on ensuring the string stability[12, 13] of the platoon.
-These systems can be classified into model-based and rule-based methods.
-Model-based approaches, such as Model Predictive Control (MPC)[14, 15, 16, 17], leverage vehicle dynamics models and optimization techniques to achieve precise control of acceleration and deceleration[18].
-However, they are often computationally demanding and may not adapt effectively to rapidly changing conditions.
-In contrast, rule-based methods rely on predefined rules and logic to dictate vehicle behavior[19, 20, 21], offering simplicity and faster computation at the cost of reduced flexibility in complex environments.
-In addition to these traditional methods, data-driven approaches, particularly those based on Reinforcement Learning (RL), have emerged as a promising alternative[13].
-RL-based CACC systems can optimize control policies by learning from real-time data[22], effectively managing complex dynamics and improving response to changing traffic conditions.
-Despite their potential, these methods require substantial training data and pose challenges in ensuring safety during the learning phase.
-CLCC enhances the maneuverability and efficiency of vehicle platoons by enabling coordinated lane changes.
-Advanced CLCC strategies, similar to CACC, include both model-based and rule-based approaches[23, 24, 25].
-Model-based methods involve predictive control techniques to optimize lane change maneuvers based on anticipated future vehicle states[26, 27], while rule-based methods utilize a set of established rules to manage lane changes according to current traffic conditions[28, 29, 30].
-Nevertheless, both approaches face limitations; model-based methods can be computationally intensive and sensitive to model inaccuracies, while rule-based methods may lack the adaptability required for managing complex traffic scenarios involving multiple interacting vehicles.
-RL-based approaches in CLCC similarly offer greater adaptability by learning from diverse traffic scenarios, allowing for optimized lane-changing strategies[31].
-While these methods enhance flexibility, they require significant data and computational resources, and ensuring reliability in mixed traffic environments remains a challenge.
-Current CACC and CLCC methods have notable shortcomings, particularly in handling scenarios with multiple types of strong interference.
-Existing strategies are typically designed to cope with one or two types of disturbances, such as sensor noise or communication delays, but struggle when faced with more complex interference patterns [32].
-This limitation can lead to safety-critical situations, especially in mixed traffic environments where unpredictable HDVs are present.
-Consequently, the current state of autonomous vehicle platooning remains relatively conservative, often prioritizing safety over efficiency to mitigate risks.
-To address these limitations, insights can be drawn from the robotics and Unmanned Aerial Vehicle (UAV) domain, where emergency obstacle avoidance and platoon reconfiguration strategies have been actively researched [33, 34, 35].
-In robotics, these strategies often employ techniques such as real-time path planning[36, 37], dynamic reconfiguration algorithms, and machine learning-based decision-making to ensure robust performance in uncertain environments.
-These methods have demonstrated success in controlled and relatively closed environments, such as factory floors or isolated test tracks.
-However, the application of these advanced strategies to autonomous vehicle platoons is limited by the open and highly variable nature of mixed traffic environments.
-Unlike controlled environments, mixed traffic flows are subject to unpredictable human behavior and diverse vehicle types, which impose significant challenges on the direct transfer of robotic strategies to autonomous driving.
-Despite these challenges, adapting and refining these approaches for use in vehicle platooning could potentially enhance the robustness and safety of autonomous driving systems.
-In light of the aforementioned challenges, this study proposes a novel approach to autonomous vehicle platooning control.
-We propose a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework.
-This framework achieves adaptive decision-making for platooning in various complex scenarios by flexibly switching between deep reinforcement learning (DRL) integrated with traffic prior knowledge and model control methods.
-Firstly, the vehicle platooning problem is modeled as a multi-agent cooperative decision-making problem, and a RL decision-making framework based on Proximal Policy Optimization (PPO)[36] is proposed, with a decision network structure and reward function designed based on traffic prior knowledge.
-A twin-world model based on safety priority judgment is introduced and coupled with the DRL algorithm to enhance algorithm safety.
-Subsequently, combining optimal control methods, a multi-scenario decision-control right adaptive switching mechanism is designed.
-To evaluate our method, we have devised intricate and high-risk mixed-traffic scenarios, as depicted in Figure 1.
-These scenarios encompass various elements such as human interference, traffic accidents, and fluctuations in traffic flow.
-Through simulations and hardware-in-the-loop experiments, our algorithm has been validated in terms of safety, robustness, and flexibility.
-Figure 1: Illustration of the considered traffic scenario and the solution. CAVs (Red) and HDVs (Green) are both considered
-Our contributions are summarized as follows:
-  • 1)
-    A Data-Model-Knowledeg Hybrid-Driven Framework for autonomous vehicle platooning is constructed, which integrates data-driven methods, model-based control, and domain-specific knowledge, enabling autonomous vehicle platoons to achieve adaptive decision-making and robust control in complex and dynamic traffic environments.
-  • 2)
-    A twin-world safety enhancement model is coupled with RL and implement an adaptive switching strategy.
-The mechanism enables the platoon to seamlessly switch between routine and safety-critical operations.
-  • 3)
-    A Cyber-Physical Experimentation Platform is established, incorporating simulation and real-world controllers, to validate the framework’s robustness and feasibility under diverse environmental conditions.
+The rapid advancements in autonomous driving and vehicle communication technologies have catalyzed significant interest in connected and autonomous vehicles (CAVs).
+Among these technological advances, autonomous vehicle platooning has emerged as a promising approach to enhance traffic efficiency, safety, and environmental sustainability [1, 2].
+By reducing inter-vehicle gaps and aerodynamic drag, vehicle platooning holds substantial potential to decrease fuel consumption and emissions[3], addressing environmental challenges and mitigating traffic congestion[4, 5].
+Current autonomous vehicle platooning systems primarily focus on two key aspects of vehicle control: Cooperative Adaptive Cruise Control (CACC) and Cooperative Lane Change Control (CLCC).
+CACC manages longitudinal dynamics to maintain inter-vehicle gaps and ensure string stability [6, 7], while CLCC facilitates lateral maneuvers for lane changes within the platoon.
+Extensive research has been dedicated to improving CACC and CLCC.
+Model-based approaches, such as Model Predictive Control (MPC), have been employed to optimize vehicle dynamics [8, 9, 10].
+However, these approaches often struggle with computational complexity and adaptability to rapidly changing environments.
+Rule-based methods offer simplicity but lack flexibility [11, 12, 13], while Reinforcement Learning (RL) has recently emerged as a promising alternative capable of learning from real-time traffic data [7, 14].
+Despite these contributions, CACC and CLCC face intrinsic limitations in managing dynamic interactions and formation reconfiguration within mixed traffic environments[15, 16], restricting their applicability to real-world scenarios characterized by complex and unpredictable conditions [17].
+Addressing these challenges necessitates a more dynamic and adaptable approach to formation maintenance and reconfiguration.
+To overcome these challenges, inspiration is drawn from formation decision-making techniques in other domains, specifically Unmanned Aerial Vehicles (UAVs).
+In the broader context of multi-agent systems, the concept of formation reconfiguration has been extensively explored in UAVs [18], which leverage dynamic formation reconfiguration to adapt to changing conditions, avoid obstacles, and fulfill mission requirements in controlled three-dimensional environments[19, 20].
+Although UAVs operate in different environments compared to ground vehicles, certain principles, such as decentralized control and dynamic adaptability, can be applied to autonomous vehicle platoons to improve their formation decision capabilities.
+Unlike UAVs, which operate in three dimensions with relatively low obstacle densities, autonomous vehicle platoons must navigate constrained road environments with frequent interactions involving human-driven vehicles (HDVs) and complex infrastructure[21].
+The unpredictable behavior of HDVs, combined with road constraints, presents a significant challenge for dynamic formation reconfiguration in vehicle platoons[22].
+Addressing these challenges requires a novel approach that integrates adaptability with the specific needs of ground-based traffic systems.
+In response, this paper proposes TriCoD, a safety-enhanced Data-Model-Knowledge Triple-Driven Adaptive Cooperative Decision-making Framework for autonomous vehicle platooning, integrating data-driven methods with model-based control to manage platoons in complex environments.
+The platooning problem is modeled as a multi-agent cooperative decision-making challenge based on Trust Region Policy Optimization (TRPO) [23], with a twin-world safety enhancement model runs in parallel with real scenarios to improve safety and adaptability during learning.
+An adaptive switching mechanism optimizes transitions between decision strategies based on traffic conditions.
+The hybrid approach is designed to enable platoon reconfiguration in complex traffic conditions.
+The system dynamically adjusts both longitudinal and lateral behaviors to enhance safety, efficiency, and flexibility in real time.
+The approach is evaluated in high-risk mixed-traffic scenarios, including human interference, incidents, and fluctuating flows, as shown in Figure 1.
+Extensive software-in-the-loop and hardware-in-the-loop experiments fully demonstrate that our framework exhibits higher security and robustness in different scenarios while taking efficiency into consideration.
+The contributions of this paper are summarized as follows:
+  • A hybrid framework for platooning decision has been developed, which integrates data-driven methods with model-based control and domain-specific knowledge, enabling platoon members to achieve adaptive decision-making in complex traffic environments.
+  • A twin-world safety enhancement model integrated with RL is proposed, which significantly enhances the sampling efficiency by providing a parallel virtual environment that allows safe exploration of diverse strategies, thereby improving the overall safety and adaptability of the system.
+  • The proposed framework uniquely enables the dissolution and reconfiguration of platoon formations in response to changing traffic conditions.
+It integrates longitudinal and lateral control, allowing the system to dynamically adjust vehicle behaviors and maintain formation integrity in complex environments.
 To better present our work, the rest of this paper is organized as follows.
-Section II presents the hybrid-driven self-organizing control framework.
-Section III introduces in detail the proposed Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework Section IV presents the performance evaluation.
-Finally, we conclude the paper in Section V.
-II Problem statement and framework review
+Section II formulates the problem in our research.
+Section III introduces in detail the proposed TriCoD framework.
+Section IV presents the performance evaluation.
+Finally, the paper is concluded in Section V.
+Figure 1: Illustration of challenging traffic conditions and our framework’s adaptive responses: formation maintenance, dissolution, and reorganization.CAVs (red) and HDVs (green) are represented with corresponding behaviors under each scenario.
+II Problem statement
 II-A Scenario Description
-As the most common application scenario for autonomous vehicle platooning, we have conducted an in-depth study on the self-organizing decision-making problem of multi-vehicle platooning on a three-lane highway with strong disturbances.
-Based on real-world issues encountered in various platooning tests, we identified and summarized several major types of strong disturbances, including human interference, traffic accidents, and traffic flow oscillations.
-In order to address these disturbances, a series of representative scenarios were devised, including emergency braking of the leading vehicle, sudden traffic accidents, and medium to low-speed traffic flow scenarios.
-These are illustrated in Figure 1.
-Conventional obstacle avoidance strategies for autonomous vehicle platooning often struggle to handle these strong disturbance factors.
-For the practical implementation of autonomous vehicle platooning, it is essential to safely navigate through regions with strong disturbances and ultimately maintain stable operation at high speeds to reach the destination.
-In this study, the high-level action decisions for the autonomous vehicle platoon are determined by our proposed Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven framework.
-This algorithm then utilizes a closed-loop Proportional-Integral-Derivative (PID) controller to translate these decisions into low-level steering and acceleration/deceleration controls.
-II-B Framework Overview
-Figure 2: The overview of the proposed data-model-knowledge hybrid-driven framework for self-organizing autonomous vehicle platooning
-To address emergency risk scenarios that traditional vehicle platooning technologies cannot handle, this study proposes a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework.
-As illustrated in Figure 2, the framework is structured into two primary layers:
-  • Upper Layer: This layer utilizes a Proximal Policy Optimization DRL decision model following the Centralized Training and Decentralized Execution (CTDE) paradigm.
-The DRL model is coupled with a twin-world model to perform safety verifications and ensure the generation of safe decisions.
-  • Lower Layer:This layer implements a Linear-Quadratic Regulator (LQR) for the longitudinal control of platoon vehicles, aimed at maintaining smaller inter-vehicle headway distances.
-These decision-making strategies are adaptively and flexibly switched using a knowledge-based finite state machine.
-The LQR controller is employed to maintain smaller headway distances in safe driving environments, thereby enhancing the platoon’s efficiency.
-In complex and challenging scenarios, the reinforcement learning model assumes control, managing the splitting and reformation of the platoon to ensure more flexible and safe driving.
+Highway environments are the most common application scenario for autonomous vehicle platooning.
+This study focuses on self-organizing decision-making for multi-vehicle platooning on a three-lane highway, addressing significant disturbances such as human interference, traffic accidents, and traffic flow oscillations, as illustrated in Figure 1.
+In these scenarios, traditional formation algorithms are no longer capable of effectively addressing the challenges posed by unpredictable disturbances, leading to reduced stability and compromised safety.
+Therefore, a new, adaptive algorithm is necessary to navigate these disturbances and ensure robust, stable, high-speed operation.
+II-B Problem Formulation
+The problem is modeled as a Partially Observable Markov Decision Process (POMDP)[24], which provides a robust framework for handling the partial observability and stochastic nature of real-world traffic scenarios.
+By employing DRL within the POMDP framework, both individual and collective goals can be balanced, enabling adaptive responses in dynamic environments.
+The system’s schematics are shown in Figure 2.
+The vehicle platoon, while making decisions as a whole, employs a decentralized execution strategy.
+The problem is formulated as a POMDP, which can be described by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{O})$, where $\mathcal{S}$ is the set of system states; $\mathcal{A}$ is the action space; $\mathcal{P}$ is the state transition model; $\mathcal{R}$ is the reward function; and $\mathcal{O}$ is the observation space.
+In this POMDP, the agent makes decisions based on the policy $\pi$ and the current observation space.
+The optimization problem is then formulated as finding the optimal policy $\pi^{*}$ that maximizes the expected sum of discounted future rewards[25]:
+$Q^{\pi}(s,a)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$ (1)
+where $\gamma\in[0,1]$ is the discount factor that balances the importance of immediate versus future rewards.
+The optimization process involves updating the policy $\pi_{\theta}$ to approximate the optimal policy $\pi^{*}$, which governs how the platoon should act to achieve long-term performance goals.
+II-B1 Observation Space $\mathcal{O}$
+In real-world highway environments, autonomous vehicles cannot directly observe the full system state due to sensor limitations and occlusions caused by other vehicles.
+Formally, the observation $o_{t}$ at time $t$ is a function of the true state $s_{t}$, but is incomplete:
+$o_{t}=f(s_{t}),\quad o_{t}\subset s_{t}$ (2)
+In this study, the observation space is structured as a matrix with dimensions $|N_{i}|\times|F|$, where $|N_{i}|$ is the number of vehicles detected by the platoon, and $|F|$ represents the feature vector describing each detected vehicle.
+The feature vector includes relative positions and velocities of other vehicles, as well as an indicator variable $\delta_{i}^{platoon}$ that specifies whether the vehicle is part of the platoon:
+$F_{i}=[x_{i},y_{i},v_{i}^{x},v_{i}^{y},\delta_{i}^{platoon}]$ (3)
+where $x_{i}$ and $y_{i}$ represent the relative longitudinal and lateral positions of vehicle $i$, and $v_{i}^{x}$ and $v_{i}^{y}$ denote its relative longitudinal and lateral velocities.
+The binary variable $\delta_{i}^{platoon}$ indicates whether vehicle $i$ is part of the platoon:
+$\delta_{i}^{platoon}=\begin{cases}1&\text{if vehicle }i\text{ belongs to the platoon},\\ 0&\text{otherwise}.\end{cases}$ (4)
+Additionally, the observation is restricted to vehicles within a sensor detection range, $D_{vision}$, ahead or behind the observing vehicle.
+Figure 2: Illustration of the system and the simulation setup.
+All high-level actions are verified through a security mask before being processed by the bottom-level controller, which will be finally converted into control signals.
+II-B2 Action Space $\mathcal{A}$
+The action space $\mathcal{A}$ defines the set of possible actions that the platoon can take at any given time step.
+In this study, each autonomous vehicle in the platoon selects from a discrete set of high-level actions, including left turn, right turn, accelerate, decelerate, and cruise.
+These actions allow the vehicles to maneuver within their lanes and adjust their speeds to maintain a safe and efficient formation.
+For a platoon consisting of $N$ autonomous vehicles, the joint action space $\mathcal{A}$ is the Cartesian product of the individual action spaces of the vehicles:
+$\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$ (5)
+Each vehicle $i$ in the platoon selects an action $a_{i}\in\mathcal{A}_{i}$, and the combined action $a_{platoon}$ is executed as:
+$a_{platoon}=(a_{1},a_{2},\ldots,a_{N})$ (6)
+In this study, a quinary encoding scheme is employed to encode the joint action space, resulting in a 125-dimensional discrete action space through the Cartesian product of the individual actions.
+After selecting high-level actions, the platoon’s low-level controllers manage steering and throttle to execute the specific vehicle movements in the simulation environment, as illustrated in Figure 2.
+II-B3 Transition Probabilities $\mathcal{P}$
+HDVs introduce randomness into the whole system, making future states inherently uncertain.
+This uncertainty is modeled using a state transition function $\mathcal{P}(s_{t+1}|s_{t},a_{t})$, which represents the stochastic nature of the environment, capturing the likelihood of transitioning from the current state $s_{t}$ to the next state $s_{t+1}$ given the current action $a_{t}$:
+$\mathcal{P}(s_{t+1}|s_{t},a_{t})=\Pr(s_{t+1}|s_{t},a_{t})$ (7)
+The transition probabilities account for the stochasticity brought by the HDVs, ensuring that the decision-making process anticipates and responds to these uncertainties.
 III Methodology
-This section details the algorithm design for autonomous vehicle platoon reconfiguration and maintenance.
-First, the problem is modeled as a Partially Observable Markov Decision Process (POMDP)[38], with a clear explanation of the observation space, action space, and reward function, integrating prior traffic knowledge.
-Then, the safety-priority-based twin-world model is introduced.
-Finally, this study combines optimal control methods to design an adaptive switching mechanism suited for multi-scenario applications, referred to as the platoon maintenance model.
-III-A Problem Formulation and Reinforcement Learning Setup
-Driving platoon vehicles on highways is a complex and challenging task, especially in mixed traffic flows.
-Treating the platoon as a single entity can lead to safety risks under strong disturbances, requiring reconfiguration.
-This process involves multi-agent cooperative decision-making[39], as each vehicle’s state is influenced by the actions of other platoon members and surrounding traffic.
-Traditional decision algorithms are inadequate for this task; thus, the problem should be modeled as a POMDP to employ deep reinforcement learning.
-The challenge lies in balancing individual and collective goals while navigating a dynamic environment where actions influence both nearby and distant vehicles.
-The system’s schematics are shown in Figure 3.
-In this study, we model the problem of multi-vehicle platooning on highways as an agent that accesses an observation at each time $t$.
-Given the sensor capabilities of platoon vehicles and the complexity of real traffic environments, the observation only provides partial information about the system state $s_{t}$.
-The platoon cannot accurately perceive the actions and intentions of all other vehicles, making the observation space $o_{t}$ a proper subset of the system state $s_{t}$.
-Using $o_{t}$, the platoon can sample decision actions $a_{t}$ from the action space according to the shared decision policy $\pi_{\theta}$, as follows:
-$a_{t}\sim\pi_{\theta}\left(a_{t}\mid o_{t}\right)$ (1)
-where the variable $\theta$ represents the trained policy parameters.
-The agents within the platoon execute action $a_{t}$, interact with the environment, and transition the system to state $s_{t+1}$, acquiring a new observation $o_{t+1}$.
-The vehicle platoon, while making decisions as a whole, employs a decentralized execution strategy.
-Therefore, we can formulate this problem as a POMDP, which can be described by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{O})$, where $\mathcal{S}$ is the set of system states, $\mathcal{A}$ is the action space, $\mathcal{P}$ is the state transition model, $\mathcal{R}$ is the reward function, and $\mathcal{O}$ is the observation space.
-The observation space $\mathcal{O}$ is determined by the system state and the position of the platoon $({o_{t}}\in{s_{t}})$.
-In the POMDP, the agent makes decisions based on the policy $\pi$ and the current observation space.
-The objective of the agent is to find the optimal policy $\pi^{*}$ that maximizes the cumulative future rewards[40].
-The optimal action-value function $Q^{\pi}(s,a)$, which can be defined as follows:
-$Q^{\pi}(s,a)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$ (2)
-The subsequent section will present the design of the observation space $\mathcal{O}$, action space $\mathcal{A}$, and reward function $\mathcal{R}$.
-  • Action space: In this study, each agent $\text{agent}_{i}$ has an action space that includes a set of high-level control decisions: left turn, right turn, accelerate, decelerate, and cruise[41, 42].
-For vehicle platooning, we integrate the action spaces of multiple agents using a quinary encoding scheme.
-This approach encodes the combined action space using a shared encoder, forming a 125-dimensional discrete action space through the Cartesian product of the individual actions.
-The overall action space for the platoon is thus the joint action space of all autonomous vehicles, denoted as $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$.
-    Figure 3: The schematics of the system and the simulation setup
-    As illustrated in Figure 3, after making high-level decisions, the platoon’s bottom-level controllers manage each vehicle’s steering and throttle, generating specific control signals to execute vehicle movements in the simulation environment.
-  • Observation Space: The observation space can be represented as a matrix with dimensions $|N_{i}|\times|F|$, where $|N_{i}|$ is the number of vehicles observed by the agent, and $|F|$ represents the feature vector of other vehicles’ positions and actions.
-The feature vector can be defined as:
-    $F_{i}=[x_{i},y_{i},v_{i}^{x},v_{i}^{y},is\_in\_platoon]$ (3)
-    To better represent the interactions between the platoon and surrounding vehicles, $x_{i},y_{i},v_{i}^{x},v_{i}^{y}$ are the relative positions and velocities of other vehicles with respect to the observing vehicle, indicating longitudinal position, lateral position, longitudinal velocity, and lateral velocity, respectively.
-The variable ${is\_in\_platoon}$ is a binary indicator (0 or 1) that denotes whether the observed vehicle is part of the platoon.
-    In this study, we consider the sensor detection range limitations and assume that only nearby vehicles can be observed.
-Specifically, only vehicles within $D_{vision}$ ahead or behind the platoon can be detected[43].
-The observation space for the entire platoon system is the Cartesian product of the individual vehicles’ observation spaces, denoted as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\cdots\times\mathcal{S}_{N}$.
-  • Reward Function: The reward function is a crucial component in training the DRL model, as it determines the model’s overall performance.
-This study designs both single-vehicle rewards and multi-vehicle collaborative rewards to ensure the robustness and safety of the platoon’s decisions.
-The overall reward for the vehicle platoon is the sum of the rewards obtained by each vehicle and the collective rewards:
-    $R_{global}(s,a)=\frac{r_{L}}{n}\sum_{i=1}^{n}R_{ind}^{i}+R_{sys}$ (4)
+This section introduces a comprehensive framework for autonomous vehicle platoon.
+The DRL-based core decision-making algorithm for formation reconfiguration is first presented to enable adaptive and secure platoon adjustments, which is followed by a formation maintenance algorithm and an adaptive switching mechanism, ensuring stable and efficient decision-making across varying traffic conditions.
+III-A Framework Overview
+Figure 3: Overview of the TriCoD framework for self-organizing autonomous vehicle platooning, which combines a data-driven upper layer and a model-driven lower layer to manage platoon adaptively.
+The upper layer uses DRL with a twin-world safety verification model for safe, real-time decision-making, while the lower layer employs a LQR for precise control in predictable conditions, ensuring stable and efficient vehicle spacing.
+Together, these layers provide robust, adaptive platooning in complex traffic environments.
+To address emergency and high-risk scenarios that traditional vehicle platooning technologies struggle to manage effectively, this study introduces TriCoD framework.
+As depicted in Figure 3, the framework seamlessly integrates two tightly coupled primary layers that collaboratively handle different aspects of platoon decision making:
+  • Upper Layer: The DRL decision model follows the Centralized Training and Decentralized Execution (CTDE) paradigm, with a twin-world safety verification model that simulates decision outcomes in a parallel virtual environment.
+This verification ensures safe, effective decision-making before deployment and enhances the DRL model’s sampling efficiency.
+The upper layer works alongside the lower layer to adapt to complex scenarios in real-time.
+  • Lower Layer: The Linear-Quadratic Regulator (LQR) provides precise control in predictable scenarios, minimizing headway for efficiency.
+This layer also collaborates with the DRL model by executing low-level control commands that align with the strategic decisions made by the upper layer.
+A knowledge-based finite state machine dynamically integrates these layers, switching decision strategies based on real-time traffic conditions.
+In stable conditions, the LQR controller optimizes efficiency with reduced headway.
+In complex scenarios, the DRL model manages platoon splitting and reformation, coordinating with the LQR for adaptive responses.
+This hybrid approach combines data-driven learning with model-based control, ensuring cohesive, adaptive responses to both routine and emergency situations.
+III-B Reward for Platooning
+The reward function is central to shaping the behavior of autonomous vehicle platoons.
+Inspired by the social behavior in AVs[26], a reward system is designed to account for both individual vehicle performance and the collective dynamics of the platoon as a whole.
+This approach ensures that each vehicle operates safely and efficiently while maintaining coordination and cohesion within the platoon.
+The total reward $R_{global}(s,a)$ is formulated as the sum of individual rewards and system-wide collaborative rewards:
+$R_{global}(s,a)=\frac{1}{n}\sum_{i=1}^{n}R_{ind}^{i}+R_{sys}$ (8)
+Both the individual vehicle perspective and the platoon-wide perspective are considered in the reward design.
+From the individual vehicle’s perspective, the reward function emphasizes safety and efficiency, focusing on collision avoidance, lane keeping, speed control, and smooth acceleration:
 $R_{ind}(s,a)=\sum_{t\in\{C,F,A\}}w_{t}\times r_{t}$ (9)
+where $r_{C}$ penalizes collisions, $r_{L}$ encourages lane discipline, $r_{F}$ rewards maintaining high speeds, and $r_{A}$ controls acceleration behavior.
+The weights $w_{C}$, $w_{L}$, $w_{F}$, and $w_{A}$ are tuned to reflect the importance of each factor, with a strong emphasis on collision avoidance through a higher value of $w_{C}$.
+From the platoon-wide perspective, the reward function promotes the cooperation and cohesion of the entire platoon.
+The system-wide reward $R_{sys}(s,a)$ encourages vehicles to remain coordinated and maintain proper formation:
 $R_{sys}(s,a)=\sum_{k\in\{M,D,H,S\}}w_{k}\times r_{k}$ (10)
-    where $w_{C}$, $w_{L}$, $w_{F}$, and $w_{A}$ are the weights for the single-vehicle rewards: collision reward $r_{C}$, lane reward $r_{L}$, high-speed reward $r_{F}$, and acceleration reward $r_{A}$, respectively.
-Since collisions are unacceptable, the absolute value of $w_{C}$ will be significantly greater than the other coefficients.
-$w_{M}$, $w_{D}$, $w_{H}$, and $w_{S}$ correspond to the multi-vehicle collaborative rewards: same-lane reward $r_{M}$, distance traveled reward $r_{D}$, headway reward $r_{H}$, and speed coordination reward $r_{S}$.
-Single-vehicle rewards primarily ensure the safety and efficiency of individual vehicles, while multi-vehicle rewards maintain the overall structure of the platoon, preventing disbandment issues.
-We considere both individual vehicle rewards and multi-vehicle collaborative rewards.
-For individual vehicle rewards, the focus is on safety and efficiency.
-In contrast, multi-vehicle collaborative rewards emphasize the cooperation and coordination among vehicles.
-The schematic diagram of the reward function is shown in Figure 8.
-    Figure 4: Illustration of the designed headway maintaining reward, which considers both the single-vehicle rewards and the multi-vehicle rewards
-III-B Deep Reinforcement Learning Algorithm
-Traditional DRL algorithms such as Deep Q-Networks (DQN) have been applied to autonomous driving decision-making[44].
-However, these algorithms primarily focus on single-agent learning and perform inadequately in scenarios requiring complex multi-vehicle cooperative decision-making.
-This study addresses the challenge of vehicle platooning reconstruction in mixed traffic flow, aiming to develop a robust self-organizing decision model for vehicle platooning under highly disruptive conditions through agent collaboration.
-To this end, we have adapted the PPO algorithm[45] to enhance its application to the safe and robust multi-vehicle cooperative decision-making problem.
-PPO is distinguished by its balance of simplicity, sample efficiency, and robustness, and has shown exceptional performance in various tasks.
-In our approach, we utilize a CTDE strategy.
-As illustrated in Figure 5, the sensor data from all platooning vehicles are aggregated into a unified observation space.
-A shared policy $\pi_{\theta}$ then outputs the actions for all vehicles within the platoon.
-Following the execution of these actions by the low-level controllers, the agents update the policy $\pi_{\theta}$ based on rewards obtained from environmental feedback, employing experience replay and gradient-based optimization.
-Figure 5: The illustration of the PPO-based formation configuration algorithm schematic model
+In this expression, $r_{M}$ rewards staying in the same lane, $r_{D}$ encourages covering greater distances, $r_{H}$ ensures proper headway between vehicles, and $r_{S}$ aligns speed coordination within the platoon.
+The corresponding weights $w_{M}$, $w_{D}$, $w_{H}$, and $w_{S}$ are designed to ensure that each vehicle’s actions contribute to the collective success of the formation.
+This reward design effectively balances the individual objectives of safety and efficiency with the collective need for coordination, ensuring that the platoon operates as a cohesive unit.
+The combined rewards guide both individual vehicles and the platoon as a whole toward achieving optimal performance in dynamic traffic conditions.
+Figure 4 illustrates the structure of the reward function, highlighting how it integrates both individual and collective factors.
+Figure 4: Illustration of the designed headway maintaining reward, which considers both the single-vehicle rewards and the multi-vehicle rewards.
+III-C Deep Reinforcement Learning Algorithm
+Traditional DRL algorithms, such as Deep Q-Networks (DQN), have been applied in autonomous driving decision-making tasks.
+However, these algorithms primarily focus on single-agent scenarios and often struggle in multi-agent cooperative decision-making environments, such as vehicle platooning.
+This study addresses the challenge of reconstructing vehicle platoons in mixed traffic, aiming to develop a robust self-organizing decision model capable of handling highly dynamic and disruptive conditions through collaborative agent behavior.
+To address this challenge, TRPO is selected for its ability to maintain stability by constraining policy updates within a trust region, thus preventing instability caused by drastic changes.
+This characteristic makes it well-suited for cooperative multi-vehicle decision-making, where stable and incremental improvements are critical.
+A CTDE strategy is employed, as shown in Figure 5, where sensor data from all platooning vehicles are aggregated into a unified observation space.
+A shared policy $\pi_{\theta}$ then generates actions for all vehicles within the platoon.
+Following the execution of these actions by the low-level controllers, the policy $\pi_{\theta}$ is updated based on environmental feedback using experience replay and gradient-based optimization.
+Figure 5: Illustration of the TRPO-based formation configuration algorithm schematic model.
 In this self-organizing autonomous vehicle platooning system, all agents are treated equally and receive the same global reward.
@@
 $\max{R_{g,t}=\sum_{i=1}^{N}r_{g,t}}$ (11)
-In deep reinforcement learning, deep neural networks (DNNs) are used to approximate policies, enabling agents to learn how to make decisions within our environment to maximize cumulative rewards.
-DNNs play a crucial role in decision-making within deep reinforcement learning.
+Deep neural networks (DNNs) are employed to approximate the policy and value functions, enabling agents to learn how to make decisions within our environment to maximize cumulative rewards.
 This study employs the Actor-Critic architecture-based DRL neural network structure depicted in Figure 6.
-Figure 6: Schematic diagram of the network structure based on the Actor-Critic architecture
+Figure 6: Illustration of the Actor-Critic Network Structure with Security-Enhanced Action Probability, which processes inputs through full conneted layers in the Actor-Critic networks.
+The Actor generates action probabilities, which are refined by a security mask to ensure safe decisions, while the Critic evaluates actions to maximize cumulative rewards.
 In this architecture, the policy loss and value function error loss are combined into a single loss function.
@@
 This study updates the network parameters using mini-batch optimization sampling. The optimization objective is to minimize the variance of Equation 15.
-III-C Twin World Security Projection
+III-D Twin World Security Projection
 While DRL enables robust decision-making in mixed traffic with strong interferences, its ”black-box” nature can lead to unexpected outcomes in standard scenarios and may not fully comply with traffic regulations.
-Although more reward functions can be used to impose these constraints, this approach can result in significant training costs and potential model divergence.
-This study addresses these issues by implementing a comprehensive and extensible safety mask.
-To ensure decision safety, we propose a safety-priority-based twin-world model.
+Although more reward functions could enforce these constraints, they may incur significant training costs and risk model divergence.
+These issues are addressed in this study through the implementation of a comprehensive and extensible safety mask.
+To ensure decision safety and sample efficiency, a safety-priority-based twin-world model is proposed.
 This model evaluates the safety of DRL decisions by first detecting potential collisions based on safety priorities, thereby reducing computational complexity.
 It then recursively computes $T_{n}$ steps in a twin world constructed from the current scenario, optimizing actions based on the platoon’s safety margin.
-Given the focus of this study on decision-making in mixed traffic, it is essential to predict the driving behavior of HDVs.
-To this end, we utilize the prevalent Intelligent Driver Model (IDM) [47] for simulating car-following behaviors.
-For lane-changing decisions, we adopt the MOBIL model [23].
-These two models, when combined, ensure a minimal risk of collisions among HDVs while closely mirroring human driving patterns.
-The safety mask models each vehicle as a circle with the vehicle’s center as the center and half the vehicle’s length as the radius.
-We simulate future $T_{n}$ steps using motion models of HDVs and CAVs to identify potential collision risks.
-While straightforward for single vehicles, this simulation becomes computationally challenging for vehicle platoons due to the exponentially increasing action space.
-To mitigate this, we propose a safety-priority-based simulation model.
-This model prioritizes vehicles with the smallest safety margins and conducts safety simulations in the twin world.
-The entire platoon tends to maintain cohesive formation, allowing for high computational efficiency suitable for real-time applications.
+Given the focus of this study on decision-making in mixed traffic, predicting the driving behavior of HDVs is essential.
+The prevalent Intelligent Driver Model (IDM) [29] is utilized for simulating car-following behaviors, while the MOBIL model [30] is adopted for lane-changing decisions.
+The combination of these models minimizes collision risk among HDVs while closely mirroring human driving patterns.
+The safety mask models each vehicle as a circle, with the center at the vehicle’s center and a radius equal to half its length.
+Future $T_{n}$ steps are simulated using motion models of HDVs and CAVs to assess collision risks.
+While this simulation is straightforward for individual vehicles, it becomes computationally challenging for platoons due to the exponentially growing action space.
+To address this, a safety-priority-based simulation model is proposed, which prioritizes vehicles with the smallest safety margins and conducts safety simulations in the twin world, ensuring efficient real-time computation and platoon cohesion.
 Safety priority is determined based on headway time, with higher priority assigned to vehicles with smaller headway times relative to HDVs. The safety priority $p_{i}$ is defined as:
@@
 Algorithm 1 Twin World Security Projection
-III-D Formation maintenance
-The platoon reconfiguration model ensures safe and robust self-organizing decision-making and control in complex scenarios with strong interference.
-However, it struggles to maintain small headway distances in routine scenarios due to overly conservative strategies aimed at avoiding collision risks.
-To address this issue, we propose an adaptive switching mechanism, the platoon maintenance model, which integrates optimal control methods.
-This model maintains stable, close headway distances in routine scenarios while dynamically adjusting driving strategies based on traffic conditions and interference levels.
-By monitoring the environment and vehicle states in real-time, the model adapts its strategies to ensure safety in complex scenarios and efficiency in routine conditions.
-This approach enhances overall coordination and driving efficiency, providing a flexible response to varying conditions.
-To address the issue of decision control allocation in vehicle platooning across different scenarios, this study employs a Finite State Machine (FSM) for decision-making.
-At any given time, the state automaton can be in only one state.
-The FSM is defined by the five-tuple $M(Q,D,\delta,q_{0},F)$, where:
+III-E Formation maintenance
+Input: $vehicles$, $riskLevel$, $Participants$, $L_{safe}$
+Output: $DecisionStrategy$
+1
+2for $i=0$ to $N$ do
+3       Assess risk level and number of traffic participants within $L_{safe}$;
+4       if ($riskLevel>=RoutineSafe$) and (Participants within $L_{safe}$) then
+5             Switch to State $S_{1}$: LQR Control Model;
+6             Activate LQR Control;
+7             Set $DecisionStrategy=LQR$;
+8
+9      else
+10             Switch to State $S_{2}$: Safety-Enhanced Data-Driven Strategy;
+11             Set $DecisionStrategy=DataDriven$;
+12
+13       end if
+14      Execute $DecisionStrategy$ based on the current state;
+15       if $DecisionStrategy==LQR$ then
+16             DecideWithLQR();
+17
+18       end if
+19      else if $DecisionStrategy==DataDriven$ then
+20             DecideWithDataDrivenStrategy();
+21
+22       end if
+23
+24 end for
+return $DecisionStrategy$
+Algorithm 2 Adaptive Switching Mechanism
+The platoon reconfiguration model ensures safe, robust decision-making in complex, high-interference scenarios but struggles with maintaining close headway distances in routine conditions due to conservative safety strategies.
+To address this, an adaptive switching mechanism, the platoon maintenance model, is proposed, incorporating optimal control methods.
+This model maintains close headway in routine scenarios while dynamically adjusting strategies based on real-time traffic conditions and interference levels, enhancing both safety in complex scenarios and efficiency in routine conditions.
+To allocate decision strategies across scenarios, a Finite State Machine (FSM) $M(Q,D,\delta,q_{0},F)$ is used, where:
   1. 1.
@@
     $D$ denotes the risk level of the traffic scenario
-This study considers a scenario as a routine safe scenario when the platoon is on the same lane and there are no other traffic participants within $L_{safe}$ ahead or behind the platoon.
-In this case, the FSM switches to state $S_{1}$, adopting the LQR control model[48] to ensure smaller headway distances.
+This study considers a scenario as safe when the platoon is on the same lane and there are no other traffic participants within $L_{safe}$ ahead or behind the platoon.
+In this case, the FSM switches to state $S_{1}$, adopting the LQR control model to ensure smaller headway distances.
 In this study, string stability within the platoon[6] is not emphasized, as it has been thoroughly addressed in previous research.
 In all other scenarios, the FSM switches to state $S_{2}$, which employs a safety-enhanced data-driven strategy to balance the safety and efficiency of the platoon.
-By defining these states and transitions, the FSM provides a structured approach to adaptively manage the control strategies of the platoon, ensuring optimal performance across varying traffic conditions.
+By defining these states and transitions, the FSM provides a structured approach to adaptively manage the decision strategies of the platoon, ensuring optimal performance across varying traffic conditions.
 The detailed procedure is illustrated in Algorithm 2.
-Input: $vehicles$, $riskLevel$, $Participants$, $L_{safe}$
-Output: $controlStrategy$
-1
-2for $i=0$ to $N$ do
-3       Assess risk level and number of traffic participants within $L_{safe}$;
-4       if ($riskLevel>=RoutineSafe$) and (Participants within $L_{safe}$) then
-5             Switch to State $S_{1}$: LQR Control Model;
-6             Activate LQR Control;
-7             Set $controlStrategy=LQR$;
-8
-9      else
-10             Switch to State $S_{2}$: Safety-Enhanced Data-Driven Strategy;
-11             Set $controlStrategy=DataDriven$;
-12
-13       end if
-14      Execute $controlStrategy$ based on the current state;
-15       if $controlStrategy==LQR$ then
-16             ControlWithLQR();
-17
-18       end if
-19      else if $controlStrategy==DataDriven$ then
-20             ControlWithDataDrivenStrategy();
-21
-22       end if
-23
-24 end for
-return $controlStrategy$
-Algorithm 2 Adaptive Switching Mechanism
 IV Simulation and Performance Evaluation
@@
 IV-A1 Environment and simulation setting
-The simulation environment is constructed upon the Highway-Env simulator[49], a reinforcement learning experimental environment library based on an OpenAI Gym environment[50].
+Figure 7: Schematic representation of the model training reward function in multiple scenarios.
+The simulation environment is constructed upon the Highway-Env simulator[31], a reinforcement learning experimental environment library based on an OpenAI Gym environment.
 As illustrated in Figure 2, all vehicles’ high-level policies are transformed into throttle and steering information through a low-level PID controller, thereby enabling vehicle movement within the simulation environment.
@@
 During the training process, a varying number of HDVs will randomly appear at randomly generated refresh points.
-The control sampling frequency is set to $15$ Hz, meaning decisions and controls are made 15 times per second.
-We use three different random seeds and trained the proposed model in various complex scenarios, with each training session running for over 100,000 steps until the model converged.
-We selecte the Deep Q-Learning (DQN)[51] and Advantage Actor-Critic (A2C)[52] algorithms as experimental baselines.
+The sampling frequency is set to $15$ Hz, meaning decisions and controls are made 15 times per second.
+Three different random seeds are used to train the proposed model in various complex scenarios, with each session running for over 100,000 steps until convergence.
+Four algorithms—DQN [32], A2C [33], PPO [27], and RecurrentPPO [34]—are selected as experimental baselines.
 The training parameters for the DRL model are detailed in Table I.
@@
 $\gamma$ Discount factor 0.8
-$w_{C}$ Weight for $r_{C}$200
-$w_{F}$ Weight for $r_{F}$ 5
-$w_{A}$ Weight for $r_{A}$ 1
-$w_{M}$ Weight for $r_{M}$ 1.2
-$w_{D}$ Weight for $r_{D}$ 1.2
-$w_{H}$ Weight for $r_{H}$ 1
-$w_{S}$ Weight for $r_{S}$ 2
 IV-A2 Performance evaluation
-In this subsection, we evaluate the performance of the proposed model within a simulated training environment, comparing it with advanced RL algorithms (e.g., DQN, A2C) and traditional CACC and CLCC platooning strategies.
-Our results indicate that the proposed vehicle platooning model significantly outperforms existing models across various scenarios.
-We designed the scenario depicted in Figure 7 not merely as a training environment but as an edge case that stresses the safety and robustness of the model under extreme conditions.
-The platoon vehicles travel at a speed of 28 m/s within a mixed traffic flow, where surrounding vehicles’ speed distribution follows a uniform range of 25 to 28 m/s.
-The traffic flow is characterized by oscillations and diverse driving styles, including aggressive, conservative, and neutral behaviors, creating a volatile environment prone to accidents.
-Furthermore, we introduced a few malfunctioning vehicles that may suddenly stall or brake unexpectedly, further heightening the risk.
-The average inter-vehicle distance for HDVs is approximately 20 meters, providing insufficient space for the platoon to execute coordinated lane changes.
-Our study trained the model in this complex and perilous scenario, incorporating randomly generated traffic flows to simulate real-world conditions.
-Figure 7: Schematic representation of the model training scenario wherein a multitude of interferences are present
-  1. 1.
-    Sample Efficiency: As shown in Figure 8, our safety-enhanced RL algorithm exhibits superior sample efficiency compared to baseline algorithms.
-The model achieves higher rewards per time step, rapidly increasing reward values within approximately 20,000 time steps and reaching initial convergence around 40,000 time steps.
-Despite typical training fluctuations, the model maintains an upward trend and stabilizes, indicating successful convergence.
-The consistently higher rewards across different scenarios are due to the model’s early prevention of illogical and illegal actions, reducing exploration time and enhancing operational efficiency.
-    TABLE II: Testing performance comparison of the collision rate, average speed and average vehicle spacing
-    of proposed model and a conventional vehicle formation model.
-    Baseline DRL Model DRL + Formation Maintenance Proposed Model
-    Collision rate(%) 61.25 5.03 5.40 3.00
-    Average Speed(km/h) 59.09 73.87 74.08 74.98
-    Average spacing(m) 5.92 5.17 3.10 3.01
-    Figure 8: Schematic representation of the model training reward function in multiple scenarios.
-  2. 2.
-    Scenario Evaluation: We focused on traffic flow oscillations, a common and representative scenario, to assess the performance of the traditional CACC and CLCC, a reinforcement learning-based model, and our proposed model.
-We analyzed the time-varying behavior of headway, acceleration, speed, and lateral position under these conditions, as illustrated in Figure 9.
-    Figure 9: Schematic illustration of headway, acceleration, velocity and lateral position of vehicle formations in three typical risk cases.
-    Safety: The traditional CACC and CLCC, influenced by the IDM-lead vehicle, displayed significant sensitivity to the motion of preceding vehicles, leading to frequent adjustments in acceleration and speed.
-This resulted in unstable behavior, particularly during traffic flow oscillations.
-In contrast, both the reinforcement learning-based model and our proposed model executed adaptive lane changes by the 7th time step, effectively mitigating the impact of oscillations and enhancing overall safety.
-    Efficiency: The traditional CACC and CLCC, lacking platoon reconfiguration capabilities, achieved efficiency gains only in the presence of large gaps.
-Concurrently, the intricate traffic patterns have precluded CLCC from identifying viable locations for collaborative lane changes, thereby impeding the potential for enhanced operational efficiency.
-Despite maintaining smaller headways, its lower speed reduced overall efficiency.
-The reinforcement learning-based model improved platoon speed by reconfiguring to more optimal lane positions.
-Our proposed model further enhanced efficiency by maintaining the platoon post-reconfiguration, sustaining high speeds and smaller headways, leading to more stable and efficient operation.
-    To generalize our findings, we tested 30,000 randomly generated high-interference scenarios to evaluate three models: the DRL model, the DRL model with formation-keeping, and our complete proposed model, including the twin-world model.
-The baseline platoon consisted of an IDM lead vehicle with LQR-following vehicles maintaining a 0.5-second headway.
-    In contrast, traditional CACC and CLCC models exhibit high collision rates in high-interference scenarios.
-However, our proposed method significantly reduces these rates while also enhancing efficiency.
-It is important to note that due to the randomness and complexity of generated scenarios, some unavoidable situations may still lead to collisions across all algorithms.
-As shown in Table II, our proposed model outperforms traditional strategies across multiple metrics.
-The DRL model effectively handles emergency scenarios, reducing the collision rate to below 5%.
-Incorporating formation-keeping maintains a comparable collision rate of 5.40%.
-However, the twin-world safety model further reduces the collision rate to 3%, underscoring the substantial impact of safety simulations.
-    Moreover, the proposed model achieves an average platoon passing speed exceeding 70 km/h, a significant improvement over the baseline’s 59.09 km/h, indicating enhanced safety and optimized traffic flow.
-Additionally, the model supports smaller headways, achieving an average spacing of 3 meters compared to 5.92 meters with traditional methods, reflecting improved vehicle control and coordination.
-    Overall, the proposed model outperforms traditional platooning strategies, demonstrating substantial improvements in safety, robustness, and efficiency.
+This section evaluates the performance of the proposed model within a simulated training environment, benchmarking it against advanced RL algorithms and conventional platooning strategies.
+Our findings indicate that the proposed platooning model consistently outperforms existing approaches across a variety of challenging scenarios.
+The scenarios, depicted in Figure 1, are designed as edge cases to test the model’s safety and robustness under extreme conditions.
+The platoon vehicles travel at a speed of 28 m/s within a mixed traffic flow, where surrounding vehicles’ speed distribution follows a uniform range of 20 to 28 m/s.
+The traffic flow is characterized by oscillations and diverse driving styles, including aggressive, conservative, and neutral behaviors, creating a volatile environment prone to accidents.Additionally, malfunctioning vehicles may stall or brake unexpectedly, further increasing risk.
+The model is trained in this complex, high-risk scenario, with randomly generated traffic flows simulating real-world conditions.
+Comparison with DRL baselines.
+Figure 7 illustrates the training performance of our model compared to baseline DRL algorithms across various scenarios.
+The training curves demonstrate that our model consistently outperforms the baseline algorithms, showing high adaptability and robustness essential for autonomous driving.
+This stability is attributed to our safety deduction model and innovative network structure, which enable our model to maintain reliable learning performance under challenging conditions and converge to a higher reward level.
+Table II provides specific performance data.
+While our model may occasionally exhibit slightly lower speed and greater average headway distance compared to some baselines, it achieves significantly higher rewards, indicating that it prioritizes safety over raw efficiency.
+In contrast, the baseline algorithms reach their performance metrics at the expense of safety, which has been further verified in simulation experiments.
+Such trade-offs are unacceptable in transportation systems where safety is paramount.
+In scenarios with high randomness and interference, such as traffic flow oscillations and traffic accidents, our model achieves significantly better convergence than other algorithms.
+This indicates an enhanced ability to learn optimal decisions under strong, dynamic disturbances, underscoring the robustness of our approach.
+TABLE II: Performance metrics comparison between proposed method and benchmarks
+Scenario Model Reward Avg. Speed Avg. HWD
+Sce-1 A2C 18.93 22.23 10.00
+DQN 9.47 25.18 20.06
+PPO 22.10 23.83 13.85
+RPPO 17.43 17.57 9.73
+Ours 23.07 20.57 9.93
+Sce-2 A2C 6.04 20.34 10.50
+DQN 2.31 27.72 9.81
+PPO 15.05 20.75 14.37
+RPPO 4.10 25.70 17.10
+Ours 31.66 20.59 11.33
+Sce-3 A2C 4.87 20.18 10.00
+DQN 11.83 20.71 12.18
+PPO 58.21 20.13 10.00
+RPPO 34.19 20.15 9.46
+Ours 61.07 20.93 8.95
+Comparison with Leading Platooning Algorithms.
+To evaluate the effectiveness of the TriCoD framework, a comparative analysis is conducted with several leading platooning algorithms.
+This analysis is based on four key performance indicators: average vehicle speed, average headway, collision rate, and pass rate.
+These metrics provide a comprehensive assessment of the model’s performance in terms of safety and efficiency.
+Notably, the pass rate metric is designed to consider scenarios where the convoy may need to halt to avoid hazards, thereby balancing risk avoidance with maintaining traffic flow.
+In the experiments, the model is evaluated against advanced lane-changing algorithms, including SuPLC[35] and SiPLC[36], as well as a baseline DRL model across different scenarios: Human Inference, Traffic Accidents, and Traffic Flow Oscillation.
+As shown in Table III, our model consistently outperformed these benchmarks in terms of collision rate and pass rate, demonstrating a unique balance between safety and operational efficiency.
+While some baseline models achieved slightly higher speeds or tighter headway distances, they do so at the expense of safety, as indicated by increased collision rates.
+In contrast, the TriCoD model maintained a zero-collision rate and a perfect pass rate of 100%, while still being able to maintain a relatively high pass rate and a smaller average headway.
+This balance illustrates our model’s capacity to prioritize safety without sacrificing operational effectiveness, especially in scenarios involving high-risk factors like traffic accidents and flow oscillations.
+Overall, these results indicate that our model consistently achieves high safety and operational metrics across varied and challenging environments, demonstrating an effective balance that baseline models do not maintain.
+TABLE III: Performance metrics comparison of multiple metrics
+between the proposed method and three state-of-the-art benchmarks
+Scenario Model Avg. Speed[m/s] Avg. HWD[m] Collision Rate Pass Rate
+Human Inference SuPLC 22.58 9.68 1.00 0.00
+SiPLC 22.58 9.68 1.00 0.00
+DRL 20.46 10.00 0.00 1.00
+Ours 20.57 9.93 0.00 1.00
+Traffic Accidents SuPLC 21.96 10.18 0.52 0.29
+SiPLC 20.81 10.15 0.54 0.21
+DRL 21.13 14.10 0.19 0.91
+Ours 20.59 11.33 0.00 1.00
+Traffic Flow Oscillation SuPLC 20.28 10.13 0.03 1.00
+SiPLC 19.68 10.15 0.03 0.97
+DRL 20.93 10.08 0.09 0.91
+Ours 20.93 8.95 0.00 1.00
 IV-B Hardware-in-loop experiment
@@
 (b) HIL experiment platform physical diagram
-Figure 10: Overall Architecture of the HIL Experimental Platform
-Figure 11: Schematic representation of case 1 results in HIL experiments
-Figure 12: Schematic representation of case 2 results in HIL experiments
+Figure 8: Illustration of the HIL platform architecture for autonomous vehicle platooning, which integrates a CARLA simulation environment, RoadRunner for road scenes and traffic flow, and domin controllers for decision-making and control.
+Figure 9: Illustration of the adaptive response to static obstruction in autonomous vehicle platooning, which decipts a scenario where the platoon encounters a multi-vehicle collision, necessitating coordinated lane changes to navigate the obstruction safely.
+Figure 10: Illustration of the adaptive response to dynamic obstruction in autonomous vehicle platooning, which depicts a scenario where the platoon leader’s sudden deceleration triggers a staggered lane-change maneuver among following vehicles to prevent collisions and maintain traffic flow.
 IV-B1 HIL experimental platform construction
-In the Hardware-in-the-Loop (HIL) experiments, we constructe an advanced simulation platform integrating a domain controller with Carla[53], a high-fidelity driving simulator, to validate our proposed model in a controlled yet realistic environment.
-The experimental highway scenario is created using Roadrunner, with varied background traffic flows generated using Highway-env.
-TABLE III: Specifications of the domain controller
+In the HIL experiments, an advanced simulation platform is constructed by integrating a domain controller with Carla [37], a high-fidelity driving simulator, to validate the proposed model in a controlled yet realistic environment.
+The experimental highway scenario is created using Roadrunner, with varied background traffic flows generated by Highway-env.
+TABLE IV: Specifications of the domain controller
 Parameter Specification
@@
 The platform consists of the domain controller, Carla simulation environment, Roadrunner, Highway-env, and PID controllers.
-The domain controller, whose parameters are shown in Table III, processes real-time traffic and road information from Carla.
+The domain controller, whose parameters are shown in Table IV, processes real-time traffic and road information from Carla.
 Carla provides detailed road conditions and vehicle dynamics, which are translated into vehicle movements via PID controllers based on high-level action commands.
@@
 This section examines the performance of our model on the HIL simulation platform in real-time autonomous platooning scenarios.
-The results indicate that our model can effectively perform perception, decision-making, and control, ensuring safe and robust autonomous platooning.
-We analyze two typical scenarios involving significant static and dynamic interferences to illustrate the model’s strategies under these conditions.
+The results indicate that our model can effectively perform perception and decision-making, ensuring safe and robust platooning.
+Two typical scenarios with significant static and dynamic interferences are analyzed to illustrate the model’s strategies.
 Demonstration videos are available on our website  [^1] Hardware-in-the-Loop Experimental Validation Video Weblink.
-In the first scenario, the platoon encounters a severe static obstruction: a multi-vehicle collision blocking multiple lanes.
-Traditional CACC strategies cannot manage coordinated deceleration due to inadequate braking distances, and CLCC strategies are hindered by vehicles behind the accident on the right.
-Our model resolves this complex situation through adaptive decision-making, avoiding secondary collisions by allowing the first vehicle to initiate a lane change and the other vehicles to follow once adequate space is available.
-As shown in Figure 11, the accident is detected at step 30.
-The first vehicle moves to the rightmost lane, completing the maneuver by step 105.
-By step 160, the remaining vehicles have sufficient space to change lanes, avoiding the collision zone, and by step 220, all vehicles have successfully moved.
-The LQR controller then stabilizes the platoon with reduced headway distances, demonstrating the model’s flexibility and efficiency in navigating static obstructions.
-In the second scenario, the platoon faces a dynamic obstruction: the leading vehicle decelerates suddenly at 5 m/s² due to a breakdown.
-Traditional strategies would struggle to avoid collisions due to short following distances and the presence of background traffic.
-Our model, however, initiates a staggered lane change, as illustrated in Figure 12.
-The first vehicle moves left at step 20, allowing the second to move right.
-The third vehicle, facing traffic in the right lane, moves left.
-By step 70, all vehicles have repositioned and resumed normal speeds.
-The rightmost lane, offering higher efficiency, sees gradual reformation of the platoon, with all vehicles settled by step 455.
+In the first scenario, the platoon encounters a multi-vehicle collision blocking several lanes.
+Traditional strategies struggle with safe deceleration and coordination, but our model enables the lead vehicle to initiate a lane change, followed by others when safe.
+As shown in Figure 9, the lead vehicle moves at step 30, with all vehicles clearing the obstruction by step 220.
+The LQR controller stabilizes the platoon with reduced headway, illustrating flexibility in static obstructions.
+In the second scenario, the lead vehicle suddenly decelerates at 5 m/s² due to a breakdown.
+Traditional models risk collision due to short distances, but ours initiates a staggered lane change (see Figure 10).
+By step 70, all vehicles are safely repositioned, and the platoon reforms in the right lane by step 455, maintaining safety and flow.
 These scenarios demonstrate the model’s capability to handle sudden obstacles by dynamically adjusting the platoon’s formation and leveraging available lane space to maintain safety and efficiency.
-The adaptive lane-changing strategy ensures minimal disruption and optimal traffic flow, showcasing the model’s robustness and flexibility in complex driving conditions.
 V Conclusion
-In the complex environment of mixed human and autonomous driving, achieving robust and safe self-organizing decision-making and control for autonomous vehicle platooning remains a significant challenge for commercial deployment.
-This paper presents a Twin-World Safety-Enhanced Data-Model-Knowledge Hybrid-Driven Adaptive Autonomous Vehicle Platooning Cooperative Control Framework for autonomous vehicle platoons.
-This framework enables adaptive decision-making in various complex scenarios by flexibly switching between reinforcement learning and model control methods, integrating traffic priors and optimal control techniques.
-Our model ensures safe and stable operation even under severe dynamic and static disturbances.
-Extensive simulations and HIL tests have demonstrated the effectiveness of the proposed model.
-The results show that our model significantly outperforms traditional platooning strategies such as CACC and CLCC in terms of safety, efficiency and robustness, demonstrating its potential for real-world deployment.
-Overall, this research contributes to the advancement of autonomous vehicle platooning by providing a model that balances safety, efficiency, and robustness.
-Future work will focus on optimizing the model’s performance, exploring its scalability to larger platoons, and conducting real-world testing to confirm its practical benefits and limitations.
+In the evolving field of autonomous driving, ensuring safe and robust self-organizing decision-making for vehicle platooning, especially in mixed traffic, presents significant challenges.
+This paper introduces TriCoD, a Twin-World Safety-Enhanced Data-Model-Knowledge Triple-Driven Adaptive Autonomous Vehicle Platooning Cooperative Decision-Making Framework.
+By integrating reinforcement learning with model control techniques, TriCoD enables adaptive decision-making across diverse, complex scenarios, guided by traffic priors and optimal control strategies.
+The proposed model significantly outperforms traditional platooning methods, such as CACC and CLCC, particularly in safety, efficiency, and robustness.
+Extensive simulations and HIL experiments validate its effectiveness in handling both static and dynamic disturbances.
+Future work will focus on optimizing the model for scalability in larger platoons, refining its performance in real-world settings, and exploring the integration of additional sensory data to enhance situational awareness.
+These advancements will further advance the reliability and applicability of autonomous platooning systems in commercial deployments.